OpenStudy (jokersmart):

The radius of a cylinder is 15 cm. The volume of the cylinder, found using 3.14 to approximate pi, is 19,782 cm3. What is the height of the cylinder? Express your answer as a whole number. cm

OpenStudy (anonymous):

The volume of cylinder is? V=πr^2h so you have volume, radius and you obviously know the constant Pi so now make h as the subject h= V/πr^2 = ???

OpenStudy (anonymous):

V = pi * r^2 * h Plug in what we know: 19782 = 3.14 * 15^2 * h Solve for h

OpenStudy (jokersmart):

Sorry for my inactiveness. I had to go eat lunch. cx is h 28??

OpenStudy (anonymous):

Yes, it is.

OpenStudy (anonymous):

:D

OpenStudy (jokersmart):

So 28 is the answer...?

OpenStudy (anonymous):

yup

OpenStudy (jokersmart):

Thank you both! 8D

OpenStudy (anonymous):

3.14 * 15^2 = 706.5 19782 / 706.5 = 28

OpenStudy (anonymous):

Np, happy to help.

OpenStudy (anonymous):

what @iGreen said is exactly the same thing as I said in my 1st post it's all about making h as the subject and know what are the values you know be confident! :)

OpenStudy (igreen):

@nopen I'm @iGreen not @iGreen.

OpenStudy (anonymous):

um.... I don't see the difference :|

OpenStudy (anonymous):

LOL

OpenStudy (igreen):

The previous person(who's still me :P) has a period..just click on me and him and you'll see that our fans and sub-scores are different.

OpenStudy (anonymous):

Oh dear...hahaha alrighty sir

OpenStudy (igreen):

Lol, I broke the fastest 99 record on that account :P 18 days